OpenStudy (anonymous):

PLEASE SOMEONE HELP !!! calculate dy/dx as a function of theta. determine the polar coordinates (r,theta), of any horizontal and vertical tangents, and convert them to Cartesian coordinates. for the function r(theta)=(4/(3-sin(theta)))

OpenStudy (anonymous):

you need to use the quotient rule on each of the parts for both dy/d(theta) and dx/d(theta). after you have done them separately you divide \[dy/d \theta/dx/d \theta \] you find that the \[(3-\sin \theta)^{2}\] drops out of the equation leaving \[\frac{ 12\cos \theta }{ -4-12\sin \theta } \] Does this help to get you on the right track?

OpenStudy (anonymous):

so i was on the right track okay. and yes it does help me re confirm what i was doing. however the main problem is how would you determine the tangents. and convert them

OpenStudy (anonymous):

would you set the top to zero and and solve. and then do the same for the bottom?

OpenStudy (anonymous):

Well, I'am not 100% about what to do next. But, I believe you need to then sub the polar values into theta, which gives some value, of dy/dx in relation to theta. (polar values being 0-2pi) once you have the value you then convert this into a Cartesian coordinate system and you should then be able to plot a tangent line

OpenStudy (anonymous):

give this website a look for some examples of this concept http://www.whitman.edu/mathematics/calculus/calculus_10_Polar_Coordinates,_Parametric_Equations_4up.pdf

OpenStudy (anonymous):

okay thanks alot im sure i can get the rest of it :) i appreciate the help

OpenStudy (anonymous):

exampe 10.7 should tell you what you need